# Memory Is Not the Bottleneck: Cost-Efficient Continual Learning via Weight Space Consolidation

### v3

@@ Abstract
Continual learning (CL) has traditionally emphasized minimizing exemplar memory usage, assuming that memory is the primary bottleneck. However, in modern computing environments—particularly those involving large foundation models—memory is inexpensive and abundant, while GPU time constitutes the main cost. This paper re-examines CL under a more realistic setting with sufficient exemplar memory, where the system can retain a representative portion of past data. We find that, under this regime, stability improves due to reduced forgetting, but plasticity diminishes as the model becomes biased toward prior tasks and struggles to adapt to new ones. Notably, even simple baselines like naive replay can match or exceed the performance of state-of-the-art methods at a fraction of the computational cost. Building on this insight, we propose a lightweight yet effective method called Weight Space Consolidation, which directly operates in the model’s weight space via two core mechanisms: (1) rank-based parameter resets to recover plasticity, and (2) weight averaging to enhance stability. Our approach outperforms strong baselines across class-incremental learning with image classifiers and continual instruction tuning with large language models, while requiring only one-third to one-fourth of the training cost. These findings challenge long-standing CL assumptions and establish a new, cost-efficient baseline for real-world continual learning systems where exemplar memory is no longer the limiting factor.
@@ 1 Introduction
As machine learning systems are increasingly considered to be deployed in dynamic, real-world environments, continual learning (CL) has emerged as a critical paradigm for adapting to evolving data streams without catastrophic forgetting [62]. A central challenge in this setting is the stability–plasticity dilemma [6, 49]: models that maintain high stability across prior tasks often fail to incorporate new knowledge (high stability, low plasticity), while those that remain highly plastic tend to forget earlier information (high plasticity, low stability). Various CL scenarios have been explored—most notably in class-incremental learning (class-IL) for image classification [47], and more recently in the continual learning of large language models (LLMs) [62, 64]. Among CL approaches, exemplar-based methods—which store and replay representative samples from past tasks—have become particularly popular due to their simplicity and effectiveness [47, 79, 64]. However, a notable trend in these methods is the use of highly constrained memory budgets. For instance, many class-IL benchmarks assume only 20 exemplars per class—roughly 4% of the total training data—are retained across tasks [54, 79]. Similarly, LLM-focused CL approaches often rely on restricted caches or memory-free mechanisms to sidestep the issue of long-term storage [64]. Yet the necessity and realism of such severe memory constraints remain questionable [10, 53, 71], and practical solutions to address this gap are still underexplored.
@@
In real-world machine learning deployments, this assumption of severely limited exemplar memory is often misaligned with practical constraints [53]. Modern storage solutions—such as cloud-based object stores or local SSDs—are both affordable and scalable. In contrast, GPU time, especially for training large-scale foundation models (e.g., LLMs), constitutes a significant bottleneck. For example, an AWS instance with 8 A100 GPUs can cost over \$30 per hour, while storing 1TB of data costs less than \$25 per month [2]. If the primary goal of CL is to enable efficient model adaptation to non-stationary data without retraining from scratch (i.e., avoiding expensive joint training), then reducing GPU cost—rather than storage usage—should be the main optimization objective.
@@
Building on this observation, we revisit the CL setting under more realistic conditions, where exemplar memory constraints are relaxed (i.e., sufficient exemplar memory is available). While prior work has largely focused on limited memory regimes, we explore how this shift impacts both algorithm design and training efficiency. To understand the implications of this regime, we analyze the stability–plasticity behavior of models trained with sufficient exemplar memory. We find that while stability improves due to reduced forgetting, plasticity diminishes as the model becomes increasingly biased toward prior tasks, resulting in difficulty adapting to new information. This observation highlights the need for mechanisms that restore plasticity without compromising stability. We further investigate how this regime affects existing CL methods across two domains: class-IL and continual instruction tuning of LLMs. In class-IL [78], we observe that many state-of-the-art methods incur significantly higher training costs—in terms of training time and VRAM usage—yet offer marginal or no improvements over naive replay baselines. In the CL of foundation models (e.g. LLMs), task-specific model merging strategies are common [15, 56, 58], but often suffer from limited plasticity or require storing a separate model per task, which limits scalability.
@@
These limitations across both domains point to the need for a new approach that is cost-efficient yet effective under sufficient memory conditions. Motivated by this, we propose Weight Space Consolidation, a simple yet effective method that operates directly in the model’s weight space. It combines (1) rank-based parameter resets, which periodically reset the dormant parameters (measured via gradient-based signal accumulation) to their pretrained values to restore plasticity, and (2) weight averaging, which maintains a running average of model weights during training to encourage convergence toward flatter, more stable optima. By modifying weights, our approach facilitates fast convergence without additional compute overhead in GPUs. Across class-IL benchmarks and LLM continual instruction tuning, our method consistently matches or surpasses the performance of state-of-the-art methods while maintaining training costs comparable to naive replay. These results demonstrate that, when exemplar memory is no longer the bottleneck, cost-efficient CL is both achievable and practical. Our contributions are summarized as follows:
@@
  • We propose a lightweight and practical method, Weight Space Consolidation, which combines rank-based parameter resets and weight averaging to address the stability–plasticity trade-off.
  • We validate our approach across class-IL tasks (e.g., CIFAR-100, ImageNet-100) and LLM continual instruction tuning (TRACE [64]), demonstrating consistent accuracy improvements and 3–4$\times$ cost reductions over state-of-the-art methods.
@@ 2 Related Works
Continual learning. Continual learning (CL) has been actively studied in various scenarios and methodological categories. Among the three scenarios of CL [60], class-incremental learning (class-IL) has been considered the most challenging and has been the most actively studied scenario [47]. Generally, CL algorithms (including class-IL) can be categorized into regularization-based approaches, which penalize changes to important parameters for past tasks [36, 1, 9, 31], rehearsal-based approaches, which store and replay exemplars from past tasks [54, 8], and expansion-based approaches, which expand the model’s capacity to balance the trade-off between stability and plasticity [69, 61]. Additional approaches focus on addressing classifier bias toward recent tasks while using the exemplars [67, 76]. While exemplar-based methods have demonstrated state-of-the-art performance, they typically rely on strict memory constraints, often limiting memory size to a small percentage of the dataset [54, 79]. Recent studies challenge the necessity of these strict memory constraints, highlighting that the computational cost of maintaining and processing memory—especially GPU usage—can far outweigh storage costs [53, 10, 22]. This shift in perspective opens the door to relaxing memory limits in order to reduce training costs, which is the focus of our work.
@@
Weight space operations. A growing body of work has explored directly manipulating model parameters in weight space across various domains, including domain generalization [66], multi-task learning [72, 70], and continual learning [46, 45, 15]. Most of these approaches operate as post hoc methods by merging the weights of pretrained models. For instance, TIES [68] proposes a selective merging strategy to mitigate interference between different tasks, while Ilharco et al. [27] demonstrate that simple arithmetic on task-specific weight deltas can edit models without further training. Building on these ideas, recent studies have extended weight-space operations to continual learning. Kozal et al. [37] apply weight averaging techniques in a CL setting, and Marczak et al. [45] introduce a selective merging approach tailored for continual adaptation. However, such methods typically require storing multiple full model checkpoints during training, fail in accumulating various task knowledge [15], and more critically, may violate the sequential constraints of CL.
In contrast, our method directly operates in weight space during training [28, 29], requiring neither multiple model copies nor post hoc merging. This enables cost-effective and online editing of the model’s parameters while maintaining compatibility with the CL setup.
@@ 3 Motivation
Notation. We generally follow the setting of continual learning (CL) [47, 79]. We consider a sequence of $T$ tasks, each associated with distribution $P_{t}$. Let $\mathcal{D}_{t}$ be the training dataset for task $t$, where $\mathcal{D}_{t}\sim P_{t}$. The tasks are presented in order $t=1,\cdots,T$. The model $F$ (its parameters $\theta$) does not retain explicit access to previous task datasets $\mathcal{D}_{j}$ for $j<t$, except via an exemplar memory buffer $\mathcal{M}$ of a capacity of $K$. Thus, at the training step of task $t$, the model updates its parameters $\theta$ using the combined data $D_{t}\cup\mathcal{M}_{1:t-1}$ and the task-designated loss $\ell(D_{t}\cup\mathcal{M}_{1:t-1};\theta)$, where $\mathcal{M}_{1:t-1}$ includes selected exemplars from earlier tasks.
@@ 3.1 Defining the sufficient exemplar memory regime
Most prior works assume a strictly limited exemplar memory budget, such that $K\ll\sum_{t=1}^{T}|\mathcal{D}_{t}|$. Under this constraint, the memory buffer $\mathcal{M}$ can retain only a small subset of examples from each past dataset $\{\mathcal{D}_{1},\cdots,\mathcal{D}_{t-1}\}$. For instance, common class-IL benchmarks typically allocate only 20 exemplars per class, which corresponds to approximately 4% of the total training data [54, 79]. As a result, the buffer provides only a partial approximation of the true task distributions $\{P_{1},\cdots,P_{t-1}\}$, leaving the model vulnerable to catastrophic forgetting [36].
By contrast, motivated by real-world scenarios where storage cost is relatively low, we define the memory buffer $\mathcal{M}$ to be sufficient if it can retain enough samples to approximate the distribution of each previous task $P_{j}$ for $1\leq j<t$. Formally, we assume the total memory budget $K$ satisfies $K\approx\kappa\sum_{j=1}^{t-1}|\mathcal{D}_{j}|$, where $\kappa\in(0,1]$ determines the fraction of past task data that can be stored. A larger value of $\kappa$ implies that $\mathcal{M}$ contains a more representative subset of earlier examples, though not necessarily the entire datasets. Under this sufficient exemplar memory setting, the mixture distribution $P_{past}$ of previously encountered tasks at the training step of task $t$ becomes:
@@
where where $\pi_{j}$ denotes the relative importance (e.g., proportional to the number of stored samples or task frequency) of each past task, and $P_{j}$ represents the corresponding data distribution. In practice, we approximate $\pi_{j}$ with its empirical counterpart $\hat{\pi}_{j}$, which can be estimated from the samples stored in the buffer. During training on task $t$, the aggregated past distribution $P_{\text{past}}$ is combined with the current task distribution $P_{t}$ to form the hybrid training distribution:
$P_{train}^{(t)}\approx\alpha_{t}P_{t}+(1-\alpha_{t})P^{(t)}_{past},$ (2)
where $\alpha_{t}\in[0,1]$ is a weighting factor that balances the influence of the current task and previously observed tasks.
@@
Figure 1: Comparison of (y-axis) average class-incremental accuracy and (x-axis) training time under different exemplar memory sizes in class-incremental learning for 10-task using CIFAR-100. As memory increases, catastrophic forgetting is mitigated (i.e., increase in accuracy), but training time (i.e., computation cost) also grows proportionally. Note that the DER, FOSTER, and MEMO are expansion-based methods (shown with X mark): FOSTER doubles the model size, while DER and MEMO scale with the number of tasks. Compared to these costly methods, Replay and Ours demonstrate high accuracy with significantly lower cost.
@@ 3.2 What changes under a sufficient exemplar memory regime?
Under sufficient exemplar memory, the distributions of previous tasks can be closely approximated, which effectively reduces forgetting—that is, improves stability. Specifically, with a large buffer $\mathcal{M}$, the empirical distributions $\tilde{P}_{j}$ of past tasks approximate their true distributions $P_{j}$ for $j<t$. This allows the empirical risk $\tilde{R}_{1:t}(\theta)$—computed over the stored exemplars—to closely approximate the ideal joint risk $R_{1:t}(\theta)=\sum_{j=1}^{t}\mathbb{E}_{x\sim P_{j}}[\ell(\theta;x)]$, as if the model were trained jointly on all tasks. As a result, the learned parameters $\tilde{\theta}^{*}_{1:t}$ remain close to the joint optimum $\theta^{*}_{1:t}$ in parameter space, preserving performance on previous tasks and mitigating catastrophic forgetting. For the complete derivation of this result, please refer to Section A.1.1.
Experimentally, Figure 1 shows that catastrophic forgetting is substantially reduced when exemplar memory is abundant. Notably, the simplest baseline (i.e., Replay) outperforms more sophisticated methods while incurring significantly lower training cost (see Figure 1). As a result, by the end of each task $t{-}1$, the model serves both as a strong minimizer for previously learned tasks $11{:}t{-}$ 1 and as a reliable initialization for the upcoming task $t$.
@@
We conjecture that as exemplar memory becomes increasingly sufficient, the model’s ability to learn new tasks (i.e., plasticity) gradually deteriorates. At task step $t$, the model is trained on a hybrid distribution $P_{\text{train}}^{(t)}\approx\alpha_{t}P_{t}+(1-\alpha_{t})P_{\text{past}}^{(t)}$ (Eq. 2), where $\alpha_{t}\in[0,1]$ controls the emphasis on the new task. As memory size increases, $\alpha_{t}$ decreases, causing the past-task distribution $P_{\text{past}}^{(t)}$ to dominate. When memory is sufficiently large, $P_{\text{past}}^{(t)}$ closely resembles $P_{\text{train}}^{(t-1)}$, and thus $P_{\text{train}}^{(t)}\approx P_{\text{train}}^{(t-1)}$. Note that this similarity results in high gradient alignment [14], which we quantify via cosine similarity $\rho_{t}=\nicefrac{{\langle\bar{g}_{t}^{(\text{new})},\bar{g}_{t}^{(\text{past})}\rangle}}{{\|\bar{g}_{t}^{(\text{new})}\|\|\bar{g}_{t}^{(\text{past})}\|}}$ (in Eq. 11), where $\bar{g}_{t}^{(\text{new})}$ and $\bar{g}_{t}^{(\text{past})}$ denote the mean gradients from the new task and past tasks, respectively. When $\rho_{t}\approx 1$, the resulting expected gradient $\bar{g}_{t}=\alpha_{t}\bar{g}_{t}^{(\text{new})}+(1-\alpha_{t})\bar{g}_{t}^{(\text{past})}$ has a reduced norm, leading to smaller parameter updates [74]:
@@
which shrinks further as $\alpha_{t}\to 0$, where $\eta$ is the step size. This limits the model’s capacity to adapt, as it tends to reuse previously learned parameters rather than learning new representations.
Such "parameter reuse" behavior has been observed when two sequential tasks are similar, leading to minimal drift across tasks [39, 13], which is a hallmark of low plasticity. Our interpretation aligns with the stability–plasticity dilemma [49, 75], where retaining prior knowledge comes at the cost of adapting to new information. It also corroborates empirical observations from [55], which showed that excessive replay can hinder the learning of new tasks under large memory conditions. Please refer to Section A.1.2 for a more detailed discussion.
Figure 2(a) compares the model’s average accuracy on new tasks across various CL methods, which is commonly used to measure the model’s ability to acquire new knowledge (i.e., plasticity) [41, 63]. Here, we experimentally confirm that as memory increases, the model’s plasticity generally degrades, resulting in lower average accuracy on new tasks. This aligns with Figure 2(b), which depicts the model’s train loss under sufficient exemplar memory. We can observe here that the large memory interferes with convergence under the CL setting. Notably in Figure 2(b), we find that reinitializing the model parameters before each task [13, 17] is a simple solution to restoring plasticity, as demonstrated by the low train loss of the model trained with reset and not continually (Orange) in Figure 2(b).
@@
(b) Train Loss under Sufficient Memory
Figure 2: Comparison of (a) average new-task accuracy and (b) training loss under different exemplar memory sizes in class-incremental learning for 10-task using CIFAR-100. As memory increases, the model’s ability to adapt to new tasks declines, resulting in reduced accuracy and slower convergence. Notably, in (b), resetting model weights before each task restores plasticity and facilitates training.

### v5

# Forget Forgetting: Continual Learning in a World of Abundant Memory

@@ Abstract
Continual learning (CL) has traditionally focused on minimizing exemplar memory, a constraint often misaligned with modern systems where GPU time, not storage, is the primary bottleneck. This paper challenges this paradigm by investigating a more realistic regime: one where memory is abundant enough to mitigate forgetting, but full retraining from scratch remains prohibitively expensive. In this practical "middle ground", we find that the core challenge shifts from stability to plasticity, as models become biased toward prior tasks and struggle to learn new ones. Conversely, improved stability allows simple replay baselines to outperform the state-of-the-art methods at a fraction of the GPU cost. To address this newly surfaced trade-off, we propose Weight Space Consolidation, a lightweight method that combines (1) rank-based parameter resets to restore plasticity with (2) weight averaging to enhance stability. Validated on both class-incremental learning with image classifiers and continual instruction tuning with large language models, our approach outperforms strong baselines while matching the low computational cost of replay, offering a scalable alternative to expensive full-retraining. These findings challenge long-standing CL assumptions and establish a new, cost-efficient baseline for real-world CL systems where exemplar memory is no longer the limiting factor.
@@ 1 Introduction
As machine learning systems are increasingly considered to be deployed in dynamic, real-world environments, continual learning (CL) has emerged as a critical paradigm for adapting to evolving data streams without catastrophic forgetting (Wang et al., 2024a). A central challenge in this setting is the stability–plasticity dilemma (Carpenter and Grossberg, 1987; Mermillod et al., 2013): models that maintain high stability across prior tasks often fail to incorporate new knowledge (high stability, low plasticity), while those that remain highly plastic tend to forget earlier information (high plasticity, low stability). Various CL scenarios have been explored—most notably in class-incremental learning (class-IL) for image classification (Masana et al., 2022), and more recently in the continual learning of large language models (LLMs) (Wang et al., 2024a; 2023). Among CL approaches, exemplar-based methods—which store and replay representative samples from past tasks—have become particularly popular due to their simplicity and effectiveness (Masana et al., 2022; Zhou et al., 2024; Wang et al., 2023). However, a notable trend in these methods is the use of highly constrained memory budgets. For instance, many class-IL benchmarks assume only 20 exemplars per class—roughly 4% of the total training data—are retained across tasks (Rebuffi et al., 2017; Zhou et al., 2024). Similarly, LLM-focused CL approaches often rely on restricted caches or memory-free mechanisms to sidestep the issue of long-term storage (Wang et al., 2023). Yet the necessity and realism of such severe memory constraints remain questionable (Chavan et al., 2023; Yousuf Harun et al., 2023), and practical solutions to address this gap are still underexplored.
@@
In real-world machine learning deployments, this assumption of severely limited exemplar memory is often misaligned with practical constraints (Prabhu et al., 2023). Modern storage solutions—such as cloud-based object stores or local SSDs—are both affordable and scalable. In contrast, GPU time, especially for training large-scale foundation models (e.g., LLMs), constitutes a significant bottleneck. For example, an AWS instance with 8 A100 GPUs can cost over \$30 per hour, while storing 1TB of data costs less than \$25 per month (Amazon Web Services, 2024). If the primary goal of CL is to enable efficient model adaptation to non-stationary data without full retraining from scratch (i.e., avoiding expensive joint training), then reducing GPU cost—rather than storage usage—should be the main optimization objective.
@@
Building on this observation, we revisit the CL setting under more realistic scenarios where exemplar memory constraints are relaxed. Our analysis reveals a critical trade-off across the memory spectrum: while traditional memory-constrained setups are often unrealistic and full-data retraining (joint training) is prohibitively expensive, a practical "middle ground" of abundant-but-not-exhaustive (i.e., sufficient) memory regime emerges. It is precisely in this realistic regime that we identify a new challenge: while stability improves due to reduced forgetting, plasticity diminishes as the model becomes biased toward prior tasks. This highlights the urgent need for cost-efficient mechanisms that can restore plasticity without sacrificing stability, a gap our work aims to fill. We further investigate how this regime affects existing CL methods across two domains: class-IL and continual instruction tuning of LLMs. In class-IL , we observe that many state-of-the-art methods incur significantly higher GPU training costs yet offer marginal improvements over naive replay baselines. In continual instruction tuning, common model merging strategies often suffer from limited plasticity or require storing a separate model per task, which limits scalability.
@@
These limitations across both domains point to the need for a new approach that is cost-efficient yet effective under abundant exemplar memory regimes. Motivated by this, we propose Weight Space Consolidation, a simple yet effective method that operates directly in the model’s weight space. It combines (1) ranking-based parameter resets, which periodically reset the dormant parameters (measured via gradient-based signal accumulation) to their pretrained values to restore plasticity, and (2) weight averaging, which maintains a running average of model weights during training to encourage convergence toward flatter, more stable optima. By modifying weights, our approach facilitates fast convergence without additional compute overhead in GPUs. Across class-IL benchmarks and LLM continual instruction tuning, our method consistently matches or surpasses the performance of state-of-the-art methods while maintaining training costs comparable to naive replay. These results demonstrate that, when exemplar memory is no longer the bottleneck, cost-efficient CL is both achievable and practical. Our contributions are summarized as follows:
@@
  • We conduct an extensive analysis across the memory spectrum to reveal the stability-plasticity trade-off, demonstrating that in the "practical" abundant memory regime, restoring lost plasticity becomes as critical as preserving stability.
  • We propose a lightweight and practical method, Weight Space Consolidation, which combines ranking-based parameter resets and weight averaging to address the stability–plasticity trade-off.
[…]
@@ 2 Related Works
Continual learning. Continual learning (CL) has been actively studied in various scenarios and methodological categories. Among the three scenarios of CL (Van de Ven and Tolias, 2019), class-incremental learning (class-IL) has been considered the most challenging and has been the most actively studied scenario (Masana et al., 2022). Generally, CL algorithms (including class-IL) can be categorized into regularization-based approaches, which penalize changes to important parameters for past tasks (Kirkpatrick et al., 2017; Aljundi et al., 2018; Cha et al., 2020; Kang et al., 2022), rehearsal-based approaches, which store and replay exemplars from past tasks (Rebuffi et al., 2017; Cha et al., 2023), and expansion-based approaches, which expand the model’s capacity to balance the trade-off between stability and plasticity (Yan et al., 2021; Wang et al., 2022). Additional approaches focus on addressing classifier bias toward recent tasks while using the exemplars (Wu et al., 2019; Zhao et al., 2020). While exemplar-based methods have demonstrated state-of-the-art performance, they typically rely on strict memory constraints, often limiting memory size to a small percentage of the dataset (Rebuffi et al., 2017; Zhou et al., 2024). Recent studies challenge the necessity of these strict memory constraints, highlighting that the computational cost of maintaining and processing memory—especially GPU usage—can far outweigh storage costs (Prabhu et al., 2023; Chavan et al., 2023; Harun et al., 2023). This shift in perspective opens the door to relaxing memory limits in order to reduce training costs, which is the focus of our work. Lastly, another line of work solely focuses on the loss of plasticity in CL (Dohare et al., 2024), where parameter resetting is commonly suggested as a solution (Ash and Adams, 2020; Galashov et al., 2024; Wang et al., 2024b; Farias and Jozefiak, 2024). In contrast, we show how both stability and plasticity are issues under realistic CL scenarios.
@@
Weight space operations. A growing body of work has explored directly manipulating model parameters in weight space across various domains, including domain generalization (Wortsman et al., 2022; Cho et al., 2025), multi-task learning (Yu et al., 2024; Yang et al., 2023), and continual learning (Marouf et al., 2025; Marczak et al., 2025; Dziadzio et al., 2024). Most of these approaches operate as post hoc methods by merging the weights of pretrained models. For instance, TIES (Yadav et al., 2024) proposes a selective merging strategy to mitigate interference between different tasks, while Ilharco et al. (Ilharco et al., 2022) demonstrate that simple arithmetic on task-specific weight deltas can edit models without further training. Building on these ideas, recent studies have extended weight-space operations to continual learning. Kozal et al. (Kozal et al., 2024) apply weight averaging techniques in a CL setting, and Marczak et al. (Marczak et al., 2025) introduce a selective merging approach tailored for continual adaptation. However, such methods typically require storing multiple full model checkpoints during training, fail in accumulating various task knowledge (Dziadzio et al., 2024), and more critically, may violate the sequential constraints of CL. In contrast, our method operates in weight space during training (Izmailov et al., 2018; Jang et al., 2025), requiring neither multiple model copies nor post hoc merging. This enables cost-effective and online editing of the model’s parameters while maintaining compatibility with the CL setup.
Positioning.
Table 1: Comparison of continual learning papers across different criteria.
Paper Constrained Memory Abundant Memory Full Memory Constrained Computation Loss of Plasticity Catastrophic Forgetting Wang et al. (2022) Prabhu et al. (2023) Harun et al. (2023) Chavan et al. (2023) Dohare et al. (2024) Galashov et al. (2024) Wang et al. (2024b) Ours
We clarify the positioning of our work in Table˜1. Several recent papers have sought cost-effective methods for CL (Prabhu et al., 2023; Harun et al., 2023; Chavan et al., 2023), but do not expand their study on varying exemplar memory scenarios (e.g., constrained/ abundant/ full). On the other hand, we enumerate our experiments across scenarios, illuminating the effect of exemplar memory sizes on the model’s stability and plasticity. While some works have already studied the loss of plasticity, they focus on extreme settings (i.e., full memory) where only plasticity is considered (Dohare et al., 2024), neglecting the issue of forgetting (Galashov et al., 2024) or failing to consider the computation costs (Wang et al., 2024b). In contrast, we focus on a realistic setting (i.e., sufficient memory) where both plasticity and stability are considered.
@@ 3 Motivation
Notation. We generally follow the setting of continual learning (CL) (Masana et al., 2022; Zhou et al., 2024). We consider a sequence of $T$ tasks, each associated with distribution $P_{t}$. Let $\mathcal{D}_{t}$ be the training dataset for task $t$, where $\mathcal{D}_{t}\sim P_{t}$. The tasks are presented in order $t=1,\cdots,T$. The model $F$ (its parameters $\theta$) does not retain explicit access to previous task datasets $\mathcal{D}_{j}$ for $j<t$, except via an exemplar memory buffer $\mathcal{M}$ of a capacity of $K$. Thus, at the training step of task $t$, the model updates its parameters $\theta$ using the combined data $D_{t}\cup\mathcal{M}_{1:t-1}$ and the task-designated loss $\ell(D_{t}\cup\mathcal{M}_{1:t-1};\theta)$, where $\mathcal{M}_{1:t-1}$ includes selected exemplars from earlier tasks.
@@ 3.1 Defining the sufficient exemplar memory regime
Most prior works assume a strictly limited exemplar memory budget, such that $K\ll\sum_{t=1}^{T}|\mathcal{D}_{t}|$. Under this constraint, the memory buffer $\mathcal{M}$ can retain only a small subset of examples from each past dataset $\{\mathcal{D}_{1},\cdots,\mathcal{D}_{t-1}\}$. For instance, common class-IL benchmarks typically allocate only 20 exemplars per class, which corresponds to approximately 4% of the total training data (Rebuffi et al., 2017; Zhou et al., 2024). As a result, the buffer provides only a partial approximation of the true task distributions $\{P_{1},\cdots,P_{t-1}\}$, leaving the model vulnerable to catastrophic forgetting.
By contrast, motivated by real-world scenarios where storage cost is relatively low but GPU cost is high, we re-examine CL in a practical regime with sufficient memory—enough to mitigate forgetting but where full retraining remains computationally expensive. Therefore, we pursue a practical “middle ground” of abundant-but-not-exhaustive exemplar memory. We define the memory buffer $\mathcal{M}$ to be sufficient if it can retain enough samples to approximate the distribution of each previous task $P_{j}$ for $1\leq j<t$. Formally, we assume the total memory budget $K$ satisfies $K\approx\kappa\sum_{j=1}^{t-1}|\mathcal{D}_{j}|$, where $\kappa\in(0,1]$ determines the fraction of past task data that can be stored. A larger value of $\kappa$ implies that $\mathcal{M}$ contains a more representative subset of earlier examples, though not necessarily the entire datasets. In Section˜3.2, we investigate when the exemplar memory size becomes sufficient by experimenting over various memory settings. Under this sufficient exemplar memory setting, the mixture distribution $P_{past}$ of previously encountered tasks at the training step of task $t$ becomes:
@@
where $\pi_{j}$ denotes the relative importance (e.g., proportional to the number of stored samples or task frequency) of each past task, and $P_{j}$ represents the corresponding data distribution. In practice, we approximate $\pi_{j}$ with its empirical counterpart $\hat{\pi}_{j}$, which can be estimated from the samples stored in the buffer. During training on task $t$, the aggregated past distribution $P_{\text{past}}$ is combined with the current task distribution $P_{t}$ to form the hybrid training distribution:
$P_{train}^{(t)}\approx\lambda P_{t}+(1-\lambda)P^{(t)}_{past}$, (2)
where $\lambda\in[0,1]$ is a factor that balances the influence of the current and previously observed tasks.
@@
Figure 1: Comparison of (y-axis) average class-incremental accuracy and (x-axis) training time under different exemplar memory sizes in class-incremental learning for 10-task using CIFAR-100. As memory increases, catastrophic forgetting is mitigated (i.e., increase in accuracy), but training time (i.e., computation cost) also grows proportionally. Note that the DER, FOSTER, and MEMO are expansion-based methods (shown with X mark): FOSTER doubles the model size, while DER and MEMO scale with the number of tasks. Compared to these costly methods, Replay and Ours demonstrate high accuracy with significantly lower cost, where our method offers the highest cost efficiency, closely approaching that of the cost lower-bound cost (i.e., Replay).
@@ 3.2 What changes under a sufficient exemplar memory regime?
Under sufficient exemplar memory, the distributions of previous tasks can be closely approximated, which effectively reduces forgetting—that is, improves stability. Specifically, with a large buffer $\mathcal{M}$, the empirical distributions $\tilde{P}_{j}$ of past tasks approximate their true distributions $P_{j}$ for $j<t$. This allows the empirical risk $\tilde{R}_{1:t}(\theta)$—computed over the stored exemplars—to closely approximate the ideal joint risk $R_{1:t}(\theta)=\sum_{j=1}^{t}\mathbb{E}_{x\sim P_{j}}[\ell(\theta;x)]$, as if the model were trained jointly on all tasks. As a result, the learned parameters $\tilde{\theta}^{*}_{1:t}$ remain close to the joint optimum $\theta^{*}_{1:t}$ in parameter space, preserving performance on previous tasks and mitigating catastrophic forgetting. For the complete derivation of this result, please refer to Section˜A.1.1.
Experimentally, Figure˜1 shows that catastrophic forgetting is substantially reduced when exemplar memory is sufficient (e.g., $|\mathcal{M}|\geq 200$). Notably, the simplest baseline (i.e., Replay) outperforms more sophisticated methods while incurring significantly lower training cost (see Figure˜1). As a result, by the end of each task t1t{-}1, the model serves both as a strong minimizer for previously learned tasks 1:t11{:}t{-}1 and as a reliable initialization for the upcoming task $t$.
@@
We find that under this condition, the challenge shifts from stability to plasticity. We conjecture that as exemplar memory becomes increasingly sufficient, the model’s ability to learn new tasks (i.e., plasticity) gradually deteriorates. At task step $t$, the model is trained on a hybrid distribution $P_{\text{train}}^{(t)}\approx\lambda P_{t}+(1-\lambda)P_{\text{past}}^{(t)}$ (Eq. 2), where $\lambda\in[0,1]$ controls the emphasis on the new task. As memory size increases, $\lambda$ decreases, causing the past-task distribution $P_{\text{past}}^{(t)}$ to dominate. When memory is sufficiently large, $P_{\text{past}}^{(t)}$ closely resembles $P_{\text{train}}^{(t-1)}$, and thus $P_{\text{train}}^{(t)}\approx P_{\text{train}}^{(t-1)}$. This similarity results in high gradient alignment (Du et al., 2018), which we quantify via cosine similarity $\rho_{t}=\nicefrac{{\langle\bar{g}_{t}^{(\text{new})},\bar{g}_{t}^{(\text{past})}\rangle}}{{\|\bar{g}_{t}^{(\text{new})}\|\|\bar{g}_{t}^{(\text{past})}\|}}$ (in Eq. 11), where $\bar{g}_{t}^{(\text{new})}$ and $\bar{g}_{t}^{(\text{past})}$ denote the mean gradients from the new tasks and past tasks, respectively. When $\rho_{t}\approx 1$, the expected gradient $\bar{g}_{t}=\lambda\bar{g}_{t}^{(\text{new})}+(1-\lambda)\bar{g}_{t}^{(\text{past})}$ has a reduced norm, causing smaller updates (Yu et al., 2020):
@@
which shrinks further as $\lambda\to 0$, where $\eta$ is the step size. This limits the model’s capacity to adapt, as it tends to reuse previously learned parameters rather than learning new representations.
Such “parameter reuse" behavior has been observed when two sequential tasks are similar, leading to minimal drift across tasks (Lee et al., 2022; Dohare et al., 2024), which is a hallmark of low plasticity. Our interpretation aligns with the stability–plasticity dilemma (Mermillod et al., 2013; Zhang et al., 2024), where retaining prior knowledge comes at the cost of adapting to new information. It also corroborates observations from Rolnick et al. (2019), which showed that excessive exemplar memory can hinder the learning of new tasks. Please refer to Section˜A.1.2 for a more detailed discussion.
Figure˜2(a) compares the model’s average accuracy on new tasks across various CL methods, which is commonly used to measure the model’s ability to acquire new knowledge (i.e., plasticity) (Liang and Li, 2023; Wang et al., 2024b). Here, we experimentally confirm that as memory increases, the model’s plasticity generally degrades, resulting in lower average accuracy on new tasks (also see Table˜6). This aligns with Figure˜2(b), which depicts the model’s train loss under full exemplar memory. We can observe here that the large memory interferes with convergence under the CL setting. Notably in Figure˜2(b), we find that reinitializing the model parameters before each task (Dohare et al., 2024; Farias and Jozefiak, 2024) is a simple solution to restoring plasticity, as demonstrated by the low train loss of the model trained with reset and not continually (Orange) in Figure˜2(b).
@@
(b) Train Loss under full exemplar memory
Figure 2: Comparison of (a) average new-task accuracy under different exemplar memory sizes and (b) training loss under full memory in class-incremental learning for 10 tasks using CIFAR-100. As memory increases, the model’s ability to adapt to new tasks declines, resulting in reduced accuracy and slower convergence. Notably, in (b), resetting model weights before each task restores plasticity and facilitates training.